OpenStudy (anonymous):

How many liters of carbon dioxide can be produced if 37.8 grams of carbon disulfide react with excess oxygen gas at 28.85 degrees Celsius and 1.02 atmospheres? CS2(l) + 3O2(g) yields CO2(g) + 2SO2(g) 2.78 liters 5.95 liters 11.9 liters 12.2 liters

OpenStudy (anonymous):

@taramgrant0543664

OpenStudy (taramgrant0543664):

Sorry about the slightly late response! Your first step is to determine the moles of carbon disulphide so you have to take the mass and divide by the molar mass

OpenStudy (taramgrant0543664):

Moles=37.8/76.14 Moles=?

OpenStudy (taramgrant0543664):

@jammy987 ?

OpenStudy (anonymous):

@taramgrant0543664 0.49

OpenStudy (anonymous):

@nincompoop

OpenStudy (taramgrant0543664):

Yes so now we can relate the moles of carbon disulfide with carbon dioxide the ratio using the stoichimometric coefficients is 1:1 so therefore 0.496 moles of carbon disulfide will make 0.496 moles of carbon dioxide

OpenStudy (taramgrant0543664):

So the next step is to plug everything into the ideal gas formula which PV=nRT where p is pressure, v is volume, n is the moles of carbon dioxide, R is the ideal gas constant (0.082) and T is temperature (make sure you change Celsius into Kalvin so add 273.15) (1.02)V=(0.496)(0.082)(302) Solve fo V

OpenStudy (taramgrant0543664):

@jammy987 sorry about the late response I was at work and I have to return to work, hopefully you get your answer!